# Classic GNNs are Strong Baselines: Reassessing GNNs for Node Classification

### v1

@@ 5.2 Influence of Hyperparameters on the Performance of GNNs
GCN 77.29 ± 0.46 73.60 ± 0.18 82.33 ± 0.19 86.33 ± 0.17
(-) Normalization 74.48 ± 1.13 71.97 ± 0.18 80.01 ± 0.48 85.21 ± 0.23
(-) Dropout 74.85 ± 0.87 72.59 ± 0.23 79.30 ± 0.37 84.47 ± 0.38
(-) Residual Connections 73.19 ± 1.46 73.01 ± 0.09 82.33 ± 0.19 79.59 ± 0.97
(-) Jumping Knowledge 77.29 ± 0.46 72.98 ± 0.24 82.33 ± 0.19 86.33 ± 0.17
GraphSAGE 82.21 ± 0.32 72.95 ± 0.31 83.89 ± 0.36 85.97 ± 0.21
(-) Normalization 77.42 ± 0.98 71.52 ± 0.17 82.12 ± 0.31 84.95 ± 0.33
(-) Dropout 80.52 ± 0.49 71.51 ± 0.04 80.36 ± 0.43 83.06 ± 0.28
(-) Residual Connections 81.75 ± 0.53 72.84 ± 0.09 83.89 ± 0.36 85.81 ± 0.45
(-) Jumping Knowledge 82.21 ± 0.32 72.76 ± 0.17 83.89 ± 0.36 85.97 ± 0.21
GAT 85.01 ± 0.46 73.30 ± 0.16 80.99 ± 0.16 86.19 ± 0.23
(-) Normalization 80.32 ± 0.83 71.68 ± 0.36 78.62 ± 0.33 84.63 ± 0.64
(-) Dropout 82.48 ± 0.34 71.83 ± 0.42 77.68 ± 0.21 85.12 ± 0.49
(-) Residual Connections 82.43 ± 0.75 72.76 ± 0.45 80.99 ± 0.16 81.37 ± 0.87
(-) Jumping Knowledge 85.01 ± 0.46 72.94 ± 0.30 80.99 ± 0.16 86.19 ± 0.23
444455556666777788889999101010109393939394949494959595959696969697979797GCN layers (homophilous graphs)Accuracy
444455556666777788889999101010107575757580808080858585859090909095959595GCN layers (heterophilous graphs)ROC-AUC/Accuracy
444455556666777788889999101010107373737373.273.273.273.273.473.473.473.473.673.673.673.673.873.873.873.8GCN layers (large-scale graph)AccuracyComputerPhotoCSRoman-EmpireMinesweeperQuestionsogbn-arxiv
4444555566667777888899991010101093939393949494949595959596969696SAGE layers (homophilous graphs)Accuracy
444455556666777788889999101010107575757580808080858585859090909095959595SAGE layers (heterophilous graphs)ROC-AUC/Accuracy
444455556666777788889999101010107272727272.572.572.572.573737373SAGE layers (large-scale graph)AccuracyComputerPhotoCSRoman-EmpireMinesweeperQuestionsogbn-arxiv
@@
Note that we do not conduct an exhaustive search regarding BN and LN, with specific choices detailed in Appendix A. We observe that the ablation of normalization does not lead to substantial deviations on small graphs. However, normalization becomes consistently crucial on large-scale graphs, where its ablation results in accuracy reductions of 3.81% and 4.69% for GraphSAGE and GAT respectively on ogbn-proteins. We believe this is because large graphs display a wider variety of node features, resulting in different data distributions across the graph. Normalization aids in standardizing these features during training, ensuring a more stable distribution.
@@
Our analysis highlights the crucial role of dropout in maintaining the performance of classic GNNs on both homophilous and heterophilous graphs, with its ablation contributing to notable accuracy declines—for instance, a 3.71% decrease for GraphSAGE on PubMed and a 5.69% decrease on Roman-Empire. This trend persists in large-scale datasets, where the ablation of dropout leads to a 2.44% and 2.53% performance decline for GCN and GAT respectively on ogbn-proteins.
@@
While the ablation of residual connections on homophilous graphs does not consistently lead to a significant performance decrease, with observed differences around 2% on Cora, Photo, and CS, the impact is more substantial on large-scale graphs such as ogbn-proteins and pokec. The effect is even more dramatic on heterophilous graphs, with the classic GNNs exhibiting the most significant accuracy reduction on Roman-Empire, for instance, a 15.72% for GCN and 12.49% for GAT. Similarly, on Minesweeper, significant performance drops were observed, emphasizing the critical importance of residual connections, particularly on heterophilous graphs. The complex structures of these graphs often necessitate deeper layers to effectively capture the diverse relationships between nodes. In such contexts, residual connections are essential for model training.
@@
As demonstrated in Figure 1, the performance trends for GCN and GraphSAGE are consistent across different graph types. On homophilous graphs and ogbn-arxiv, both models achieve optimal performance with a range of 4 to 8 layers. In contrast, on heterophilous graphs, their performance improves with an increasing number of layers, indicating that deeper networks are more beneficial for these graphs. We discuss scenarios with more than 10 layers in Appendix B.
Observation 5: Jumping knowledge mode has minimal influence on large-scale graphs.
The ablation of this component typically results in marginal accuracy decreases of less than 0.5% for classic GNNs across large-scale graphs, suggesting the lesser impact of jumping knowledge than other components. By comparison, on smaller-scale homophilous graphs, such as CiteSeer and CS, the ablation of jumping knowledge can lead to decreases of around 1.5 to 2% in the model performance. On some heterphilous graphs, the effect is also significant, with the accuracy of GAT decreasing by 2.32% on Chameleon, GraphSAGE by 3.15% and GAT by 3.84% on Minesweeper.
@@ A.2 Hyperparameters and Reproducibility
For the hyperparameter selections of classic GNNs, in addition to what we have covered, we list other settings in Tables 8, 9, 10. Notably, for heterophilous graphs, we expand the search range for the number of layers to include three additional settings: $\{11,12,15,20\}$ (See Sec. B.2 for further analysis). This adjustment is based on our empirical evidence suggesting that deep networks tend to yield performance improvements on heterophilous graphs. The ReLU function serves as the non-linear activation. Further details regarding hyperparameters can be found in our code https://github.com/LUOyk1999/tunedGNN.
Due to the large size of the graphs in ogbn-proteins, ogbn-products, and pokec, which prevents full-batch training on GPU memory, we adopt different training strategies. For ogbn-proteins, we utilize the optimized neighbor sampling method [20]. For pokec and ogbn-products, we apply the random partitioning method previously used by GTs [12, 71, 70] to enable mini-batch training. For other datasets, we employ full-batch training. In all experiments, we use the validation set to select the best hyperparameters. GNN denotes our implementation of the GNN model.
@@
Dataset JK ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora False True False 0.5 3 256 0.005 500
Citeseer False True False 0.5 3 256 0.005 500
Pubmed False True False 0.5 2 64 0.01 500
Computer True False LN 0.7 4 512 0.001 1000
Photo True True LN 0.0 9 512 0.001 1000
CS True True LN 0.3 7 512 0.001 1500
Physics True True LN 0.5 9 512 0.001 1500
WikiCS True True False 0.5 10 512 0.001 1000
Squirrel True False BN 0.5 8 64 0.01 500
Chameleon True True BN 0.5 5 256 0.01 200
Amazon-Ratings True True False 0.3 10 512 0.001 2500
Roman-Empire True True False 0.3 11 512 0.001 2500
Minesweeper True True False 0.3 15 512 0.001 2000
Questions True True False 0.2 11 512 0.001 1500
ogbn-proteins False True BN 0.3 3 512 0.01 100
ogbn-arxiv True True BN 0.5 6 256 0.0005 2000
ogbn-products False False LN 0.5 5 256 0.003 300
pokec False True BN 0.2 7 256 0.0005 2000
@@
Dataset JK ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora False True False 0.5 6 256 0.005 500
Citeseer True False False 0.5 5 256 0.001 500
Pubmed False False False 0.5 3 256 0.001 500
Computer True True LN 0.7 7 512 0.001 1000
Photo True True False 0.7 5 512 0.001 1000
CS True True LN 0.0 7 512 0.001 1500
Physics True True LN 0.5 4 512 0.001 1500
WikiCS True True LN 0.5 5 512 0.001 1000
Squirrel True True False 0.0 2 64 0.005 500
Chameleon False True False 0.5 4 256 0.001 200
Amazon-Ratings True True LN 0.3 5 512 0.001 2500
Roman-Empire True True False 0.3 11 512 0.001 2500
Minesweeper True True False 0.3 11 512 0.001 2000
Questions True True LN 0.0 11 512 0.001 1500
ogbn-proteins False True BN 0.3 6 512 0.01 1000
ogbn-arxiv True True BN 0.5 7 256 0.0005 2000
ogbn-products False False LN 0.5 5 256 0.003 1000
pokec False True BN 0.2 7 256 0.0005 2000
@@
Dataset JK ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora True True False 0.5 6 256 0.001 500
Citeseer True False False 0.5 4 256 0.001 500
Pubmed False False False 0.5 2 256 0.001 500
Computer True True LN 0.7 6 512 0.001 1000
Photo True True LN 0.0 5 512 0.001 1000
CS True True LN 0.3 5 512 0.001 1500
Physics True True LN 0.0 5 512 0.001 1500
WikiCS True True False 0.5 7 512 0.001 1000
Squirrel False False False 0.0 3 256 0.001 500
Chameleon True True BN 0.5 4 256 0.005 200
Amazon-Ratings True True False 0.3 7 512 0.001 2500
Roman-Empire False True LN 0.3 6 512 0.001 2500
Minesweeper True True False 0.3 15 512 0.001 2000
Questions False True LN 0.2 4 512 0.001 1500
ogbn-proteins False True BN 0.3 7 512 0.01 1000
ogbn-arxiv True True BN 0.5 6 256 0.0005 2000
ogbn-products False False LN 0.5 5 256 0.003 1000
pokec False True BN 0.2 7 256 0.0005 2000
@@
Appendix B Additional Benchmark Results
B.1 GAT with Edge Features on ogbn-proteins
While DeepGCN [33] introduced training models up to 56 layers deep and DeeperGCN [34] further extended this to 112 layers, our experiments suggest that such depth is not necessary. Specifically, while the DeeperGCN achieved an accuracy of 85.50% on ogbn-proteins, it utilized edge features as input, a configuration not commonly employed in the standard baselines of the OGB dataset [24]. As our experiments do not incorporate edge features on ogbn-proteins, we exclude DeeperGCN from our main text to maintain a fair comparison.
@@ B.2 Deeper Networks on Heterophilous Graphs
On heterophilous graphs, the performance of classic GNNs improves with an increasing number of layers limited to 10, as evidenced by Figure 1 in the main text. We explore scenarios with more than 10 layers in this subsection. Specifically, we consider GCN and GAT with layer configurations of 11, 12, 15, and 20 for the Amazon-Ratings, Roman-Empire, and Minesweeper datasets. The results are shown in Table 12. The variation in the optimal number of layers ($L$) could stem from the distinct structures inherent in different graphs. Heterophilous graphs may have more complex structures, thus necessitating a higher $L$. However, the slight improvements observed with larger $L$ values suggest that very deep networks may not yield significantly better results. Overall, the best results for classic GNNs are achieved when $L$ is limited to 15.
Amazon-Ratings Roman-Empire Minesweeper
[…]
GCN ($L=11$) 53.52 ± 0.19 91.35 ± 0.37 97.01 ± 0.18
GCN ($L=12$) 53.66 ± 0.11 90.75 ± 0.37 97.16 ± 0.28
GCN ($L=15$) 53.65 ± 1.02 91.10 ± 0.41 97.77 ± 0.38
GCN ($L=20$) 53.50 ± 0.58 91.10 ± 0.76 97.76 ± 0.30
GAT ($L=11$) 54.61 ± 0.21 91.07 ± 0.40 97.11 ± 0.09
GAT ($L=12$) 54.55 ± 0.18 90.79 ± 0.44 97.59 ± 0.15
GAT ($L=15$) 55.01 ± 0.45 90.98 ± 0.41 97.76 ± 0.37
GAT ($L=20$) 55.01 ± 0.17 91.30 ± 0.38 97.56 ± 0.47
Appendix C Limitations & Broader Impacts

### v2

@@ 5.2 Influence of Hyperparameters on the Performance of GNNs
GCN 77.29 ± 0.46 73.53 ± 0.12 82.33 ± 0.19 86.33 ± 0.17
(-) Normalization 74.48 ± 1.13 71.53 ± 0.14 80.01 ± 0.48 85.21 ± 0.23
(-) Dropout 74.85 ± 0.87 72.06 ± 0.13 79.30 ± 0.37 84.47 ± 0.38
(-) Residual Connections 73.19 ± 1.46 72.91 ± 0.17 - 79.59 ± 0.97
GraphSAGE 82.21 ± 0.32 73.00 ± 0.28 83.89 ± 0.36 85.97 ± 0.21
(-) Normalization 77.42 ± 0.98 71.13 ± 0.27 82.12 ± 0.31 84.95 ± 0.33
(-) Dropout 80.52 ± 0.49 71.30 ± 0.21 80.36 ± 0.43 83.06 ± 0.28
(-) Residual Connections 81.75 ± 0.53 72.22 ± 0.49 - 85.81 ± 0.45
GAT 85.01 ± 0.46 73.30 ± 0.18 80.99 ± 0.16 86.19 ± 0.23
(-) Normalization 80.32 ± 0.83 71.33 ± 0.29 78.62 ± 0.33 84.63 ± 0.64
(-) Dropout 82.48 ± 0.34 71.68 ± 0.32 77.68 ± 0.21 85.12 ± 0.49
(-) Residual Connections 82.43 ± 0.75 72.47 ± 0.34 - 81.37 ± 0.87
222233334444555566667777888899991010101093939393949494949595959596969696GCN layers (homophilous graphs)Accuracy
22223333444455556666777788889999101010107575757580808080858585859090909095959595GCN layers (heterophilous graphs)ROC-AUC/Accuracy
22223333444455556666777788889999101010107272727272.572.572.572.57373737373.573.573.573.5GCN layers (large-scale graph)AccuracyComputerPhotoCSRoman-EmpireMinesweeperQuestionsogbn-arxiv
22223333444455556666777788889999101010109393939394949494959595959696969697979797SAGE layers (homophilous graphs)Accuracy
22223333444455556666777788889999101010107575757580808080858585859090909095959595SAGE layers (heterophilous graphs)ROC-AUC/Accuracy
222233334444555566667777888899991010101071.571.571.571.57272727272.572.572.572.573737373SAGE layers (large-scale graph)AccuracyComputerPhotoCSRoman-EmpireMinesweeperQuestionsogbn-arxiv
@@
We observe that the ablation of normalization does not lead to substantial deviations on small graphs. However, normalization becomes consistently crucial on large-scale graphs, where its ablation results in accuracy reductions of 4.79% and 4.69% for GraphSAGE and GAT respectively on ogbn-proteins. We believe this is because large graphs display a wider variety of node features, resulting in different data distributions across the graph. Normalization aids in standardizing these features during training, ensuring a more stable distribution.
@@
Our analysis highlights the crucial role of dropout in maintaining the performance of classic GNNs on both homophilous and heterophilous graphs, with its ablation contributing to notable accuracy declines—for instance, a 2.70% decrease for GraphSAGE on PubMed and a 6.57% decrease on Roman-Empire. This trend persists in large-scale datasets, where the ablation of dropout leads to a 2.44% and 2.53% performance decline for GCN and GAT respectively on ogbn-proteins.
@@
While the ablation of residual connections on homophilous graphs does not consistently lead to a significant performance decrease, with observed differences around 2% on Cora, Photo, and CS, the impact is more substantial on large-scale graphs such as ogbn-proteins and pokec. The effect is even more dramatic on heterophilous graphs, with the classic GNNs exhibiting the most significant accuracy reduction on Roman-Empire, for instance, a 16.43% for GCN and 5.48% for GAT. Similarly, on Minesweeper, significant performance drops were observed, emphasizing the critical importance of residual connections, particularly on heterophilous graphs. The complex structures of these graphs often necessitate deeper layers to effectively capture the diverse relationships between nodes. In such contexts, residual connections are essential for model training.
@@
As demonstrated in Figure 1, the performance trends for GCN and GraphSAGE are consistent across different graph types. On homophilous graphs and ogbn-arxiv, both models achieve optimal performance with a range of 2 to 6 layers. In contrast, on heterophilous graphs, their performance improves with an increasing number of layers, indicating that deeper networks are more beneficial for these graphs. We discuss scenarios with more than 10 layers in Appendix B.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
We extend our gratitude to Yiwen Sun for her invaluable assistance. We also express our appreciation to all the anonymous reviewers and ACs for their insightful and constructive feedback. This work received support from National Key R&D Program of China (2021YFB3500700), NSFC Grant 62172026, National Social Science Fund of China 22&ZD153, the Fundamental Research Funds for the Central Universities, State Key Laboratory of Complex & Critical Software Environment (CCSE), HK PolyU Grant P0051029, HK PolyU Grant P0038850, and HK ITF Grant ITS/359/21FP. Lei Shi is with Beihang University and State Key Laboratory of Complex & Critical Software Environment.
@@ A.2 Hyperparameters and Reproducibility
For the hyperparameter selections of classic GNNs, in addition to what we have covered, we list other settings in Tables 8, B.1, B.1. Notably, for heterophilous graphs, we expand the search range for the number of layers to include three additional settings: $\{12,15,20\}$ (See Section B.2 for further analysis). This adjustment is based on our empirical evidence suggesting that deep networks tend to yield performance improvements on heterophilous graphs. The ReLU function serves as the non-linear activation. Further details regarding hyperparameters can be found in our code https://github.com/LUOyk1999/tunedGNN.
Due to the large size of the graphs in ogbn-proteins, ogbn-products, and pokec, which prevents full-batch training on GPU memory, we adopt different batch training strategies. For ogbn-proteins, we utilize the optimized neighbor sampling method [20]. For pokec and ogbn-products, we apply the random partitioning method previously used by GTs [12, 76, 75] to enable mini-batch training. For other datasets, we employ full-batch training. In all experiments, we use the validation set to select the best hyperparameters.
The testing accuracy achieved by the model that reports the highest result on the validation set is used for evaluation. Additionally, we report mean scores and standard deviations after 5 independent runs with different initializations.
@@
Dataset ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora False False 0.7 3 512 0.001 500
Citeseer False False 0.5 2 512 0.001 500
Pubmed False False 0.7 2 256 0.005 500
Computer False LN 0.5 3 512 0.001 1000
Photo True LN 0.5 6 256 0.001 1000
CS True LN 0.3 2 512 0.001 1500
Physics True LN 0.3 2 64 0.001 1500
WikiCS False LN 0.5 3 256 0.001 1000
Squirrel True BN 0.7 4 256 0.01 500
Chameleon False False 0.2 5 512 0.005 200
Amazon-Ratings True BN 0.5 4 512 0.001 2500
Roman-Empire True BN 0.5 9 512 0.001 2500
Minesweeper True BN 0.2 12 64 0.01 2000
Questions True False 0.3 10 512 0.001 1500
ogbn-proteins True BN 0.3 3 512 0.01 100
ogbn-arxiv True BN 0.5 5 512 0.0005 2000
ogbn-products False LN 0.5 5 256 0.003 300
pokec True BN 0.2 7 256 0.0005 2000
Appendix B Additional Benchmarking Results
B.1 GAT with Edge Features on ogbn-proteins
While DeepGCN [33] introduced training models up to 56 layers deep and DeeperGCN [34] further extended this to 112 layers, our experiments suggest that such depth is not necessary. Specifically, while the DeeperGCN achieved an accuracy of 85.50% on ogbn-proteins, it utilized edge features as input, a configuration not commonly employed in the standard baselines of the OGB dataset [24]. As our experiments do not incorporate edge features on ogbn-proteins, we exclude DeeperGCN from the main text to maintain a fair comparison.
@@ B.1 GAT with Edge Features on ogbn-proteins
Dataset ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora False False 0.7 3 256 0.001 500
Citeseer False False 0.2 3 512 0.001 500
Pubmed False False 0.7 4 512 0.005 500
Computer False LN 0.3 4 64 0.001 1000
Photo True LN 0.2 6 64 0.001 1000
CS True LN 0.5 2 512 0.001 1500
Physics True BN 0.7 2 64 0.001 1500
WikiCS False LN 0.7 2 256 0.001 1000
Squirrel True BN 0.7 3 256 0.01 500
Chameleon True BN 0.7 4 256 0.01 200
Amazon-Ratings True BN 0.5 9 512 0.001 2500
Roman-Empire False BN 0.3 9 256 0.001 2500
Minesweeper True BN 0.2 15 64 0.01 2000
Questions False LN 0.2 6 512 0.001 1500
ogbn-proteins True BN 0.3 6 512 0.01 1000
ogbn-arxiv True BN 0.5 4 256 0.0005 2000
ogbn-products False LN 0.5 5 256 0.003 1000
pokec True BN 0.2 7 256 0.0005 2000
@@
Dataset ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora True False 0.2 3 512 0.001 500
Citeseer True False 0.5 3 256 0.001 500
Pubmed False False 0.5 2 512 0.01 500
Computer False LN 0.5 2 64 0.001 1000
Photo True LN 0.5 3 64 0.001 1000
CS True LN 0.3 1 256 0.001 1500
Physics True BN 0.7 2 256 0.001 1500
WikiCS True LN 0.7 2 512 0.001 1000
Squirrel True BN 0.5 7 512 0.005 500
Chameleon True BN 0.7 2 256 0.01 200
Amazon-Ratings True BN 0.5 4 512 0.001 2500
Roman-Empire True BN 0.3 10 512 0.001 2500
Minesweeper True BN 0.2 15 64 0.01 2000
Questions True LN 0.2 3 512 0.001 1500
ogbn-proteins True BN 0.3 7 512 0.01 1000
ogbn-arxiv True BN 0.5 5 256 0.0005 2000
ogbn-products False LN 0.5 5 256 0.003 1000
pokec True BN 0.2 7 256 0.0005 2000
@@
Roman-Empire Minesweeper
[…]
GCN ($L=12$) 90.68 ± 0.44 97.76 ± 0.24
GCN ($L=15$) 90.74 ± 0.38 97.65 ± 0.81
GCN($L=20$) 90.43 ± 0.52 97.52 ± 0.28
GrapSAGE ($L=12$) 90.96 ± 0.46 97.02 ± 0.51
GrapSAGE ($L=15$) 90.78 ± 0.63 97.77 ± 0.62
GrapSAGE ($L=20$) 90.22 ± 0.69 97.73 ± 0.88
@@
Table 13: Node classification results over homophilous graphs (%). + indicates the implementation of classic GNNs using JK as a hyperparameter configuration in our past experiments.
Cora CiteSeer PubMed Computer Photo CS Physics WikiCS
[…]
GCN+ 85.08 ± 0.52 72.98 ± 0.84 81.32 ± 0.72 93.80 ± 0.29 96.51 ± 0.20 95.80 ± 0.28 97.43 ± 0.05 80.27 ± 0.71
GraphSAGE+ 84.18 ± 0.81 71.93 ± 0.85 79.41 ± 0.53 93.59 ± 0.22 96.41 ± 0.17 96.12 ± 0.24 97.21 ± 0.05 80.51 ± 0.48
GAT+ 84.64 ± 1.27 72.10 ± 1.10 79.70 ± 0.70 93.93 ± 0.16 96.67 ± 0.13 96.08 ± 0.10 97.30 ± 0.06 80.75 ± 0.74
Table 14: Node classification results on heterophilous graphs (%).
[…]
GCN+ 44.50 ± 1.92 46.11 ± 3.16 53.57 ± 0.32 91.35 ± 0.37 97.77 ± 0.38 77.40 ± 1.07
GraphSAGE+ 39.93 ± 1.58 43.44 ± 3.19 54.72 ± 0.38 92.19 ± 0.58 96.95 ± 0.41 77.96 ± 0.72
GAT+ 38.72 ± 1.46 43.44 ± 3.00 54.99 ± 0.71 91.60 ± 0.21 97.76 ± 0.37 79.04 ± 1.27
Table 15: Node classification results on large-scale graphs (%).
ogbn-proteins ogbn-arxiv ogbn-products pokec
[…]
GCN+ 77.29 ± 0.46 73.60 ± 0.18 82.33 ± 0.19 86.33 ± 0.17
GraphSAGE+ 82.21 ± 0.32 72.95 ± 0.31 83.89 ± 0.36 85.97 ± 0.21
GAT+ 85.01 ± 0.46 73.30 ± 0.16 80.99 ± 0.16 86.19 ± 0.23
@@ B.2 Deeper Networks on Heterophilous Graphs
On heterophilous graphs, the performance of classic GNNs improves with an increasing number of layers limited to 10, as evidenced by Figure 1 in the main text. We explore scenarios with more than 10 layers in this subsection. Specifically, we consider GCN and GraphSAGE with layer configurations of 12, 15, and 20 for the Roman-Empire and Minesweeper datasets. The results are shown in Table 12. The variation in the optimal number of layers ($L$) could stem from the distinct structures inherent in different graphs. Heterophilous graphs may have more complex structures, thus necessitating a higher $L$. However, the slight improvements observed with larger $L$ values suggest that very deep networks may not yield significantly better results. Overall, the best results for classic GNNs are achieved when $L$ is limited to 15.
B.3 Jumping Knowledge Mode and Early Results
where $L$ is the number of GNN layers. In our previous experimental setups, we treated JK as a hyperparameter configuration for GNNs. Based on the hyperparameter configurations outlined in Section 3, we expanded the tuning space to include the decision of whether to use JK. In past experiments, we did not perform an exhaustive search; instead, we selected subsets based on experience within this search space, and our early results are reported in Table B.1, B.1, and B.1 (For additional information, please refer to https://arxiv.org/abs/2406.08993v1). However, after a more detailed hyperparameter tuning, we found that JK may not be necessary. In most datasets, the results without using JK are comparable to, and sometimes even better than, those with JK. Consequently, we removed JK from the hyperparameter tuning search space in our paper.
Appendix C Visualization
Here, we present t-SNE visualizations of classification results. As shown in Figure LABEL:fig:2, the node embeddings generated by GCN (our implementation) display greater inter-class distances than those produced by Polynormer.
Appendix D Limitations & Broader Impacts